OpenStudy (anonymous):

(1,-8) m=-6 point-slope into slope-intercept how do i do this

OpenStudy (anonymous):

y - y1 = m*(x-x1) plug in

OpenStudy (anonymous):

yeah i have already done that and basically did all the math but i cant figure out how to put it into slope intercept

OpenStudy (anonymous):

ok plug in the number and show me and i will help u

OpenStudy (anonymous):

y+8=-6(x-1)

OpenStudy (anonymous):

bow distribute the -6

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

y+8=-6x+6

OpenStudy (anonymous):

now, how can you move the 8 to the other side?

OpenStudy (anonymous):

by subtracting it?

OpenStudy (anonymous):

correct do that

OpenStudy (anonymous):

y=-6x+-2 ?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

see how easy it was, and u did all the work

OpenStudy (anonymous):

yeah i had already done the work but i was getting confused about something, anyways thanks a bunch :)